OpenStudy (anonymous):

Solve. 9n + 7 + 3n = 103

OpenStudy (anonymous):

n=8

OpenStudy (anonymous):

9n +3n = 103 - 7 12n = 96 n = 96/12 n = 8

OpenStudy (anonymous):

Thankk Youu !

OpenStudy (anonymous):

add 9n and 3n 12n+7=103 12n=96 n=8